# Deep Learning, Machine Learning - Digital Signal and Image Processing: From Theory to Application

### v1

@@ 11.2.1 Huffman Coding
Huffman coding is a widely used lossless compression algorithm that assigns variable-length codes to symbols (in this case, pixel values), with shorter codes assigned to more frequently occurring symbols. This is based on the observation that, in most images, some pixel values occur more frequently than others. By encoding these frequent values with fewer bits and less frequent values with more bits, Huffman coding achieves compression [128].
@@ 11.2.2 Run-length Encoding (RLE)
Run-length encoding (RLE) is a simple lossless compression technique that is particularly effective for images with large contiguous regions of the same pixel value, such as binary or grayscale images. The basic idea is to represent consecutive occurrences of the same value as a pair of the value and the count of how many times it repeats [33].
@@ 11.4 Lossy Compression Techniques
Lossy compression techniques are widely used in digital image and multimedia storage because they significantly reduce file sizes by discarding less critical information. Unlike lossless compression, which preserves every detail, lossy compression allows some degradation of image quality in exchange for better compression ratios. The key to effective lossy compression is to remove information that is perceptually less important to the human visual system. In this section, we will cover two of the most important techniques used in lossy compression: the Discrete Cosine Transform (DCT) and the Discrete Wavelet Transform (DWT). These methods are foundational in many compression standards, such as JPEG [89].
@@ 11.4.1 Discrete Cosine Transform (DCT)
The Discrete Cosine Transform (DCT) is a fundamental tool in image compression. It converts an image from the spatial domain, where pixel intensities are represented, into the frequency domain, where the image is represented as a sum of cosine functions at various frequencies. This transformation allows the separation of the image into low-frequency components (which represent the bulk of the image’s structure) and high-frequency components (which represent fine details and edges) [4].
@@
  • Advantages: The DCT provides good energy compaction and can be efficiently implemented. It is particularly effective for smooth images with few sharp edges [158].
  • Limitations: The DCT assumes that the image is locally smooth, which means it may not perform as well for images with many sharp edges or textures. Additionally, blocking artifacts can occur if the image is divided into small blocks for compression (as in JPEG) [158].
@@ 11.4.2 Discrete Wavelet Transform (DWT)
The Discrete Wavelet Transform (DWT) is another powerful tool used in lossy image compression. Unlike the DCT, which operates on blocks of fixed size, the DWT provides a multi-resolution representation of the image by decomposing it into different levels of detail. The DWT captures both frequency and spatial information, which makes it well-suited for images with varying levels of detail, such as natural scenes [220].
@@
The DWT decomposes an image into a set of subbands, each corresponding to different frequency components and spatial resolutions [172]. Typically, the image is first decomposed into four subbands:
@@ 9 LL, (LH, HL, HH) = coeffs2
14 axes[0, 0].imshow(LL, cmap=’gray’)
15 axes[0, 0].set_title(’Approximation (LL)’)
@@
23 axes[1, 1].imshow(HH, cmap=’gray’)
24 axes[1, 1].set_title(’Diagonal Detail (HH)’)
@@
  • Advantages: The DWT provides a multi-resolution representation of the image, which allows for efficient compression by discarding or heavily quantizing high-frequency components. It is especially effective for images with detailed textures and varying levels of detail [209].
  • Limitations: Wavelet-based compression can be more computationally intensive than DCT-based methods. Additionally, wavelet compression can introduce ringing artifacts around sharp edges if the high-frequency components are too aggressively quantized [209].
@@ 11.5 JPEG Compression Standard
JPEG (Joint Photographic Experts Group) is one of the most commonly used image compression standards, known for its ability to achieve high compression ratios with minimal perceptual loss in image quality. The JPEG algorithm is based on the Discrete Cosine Transform (DCT) and employs several key steps to compress images in a lossy fashion [204]. In this section, we will break down the main steps of the JPEG compression process and explain its widespread use and performance for still image compression.
@@ 11.5.2 Widespread Use and Performance of JPEG
JPEG is widely used for still image compression because it offers a good balance between compression ratio and image quality. It is particularly effective for natural images (e.g., photographs) where slight loss of detail is acceptable. JPEG can achieve compression ratios of 10:1 or higher with little perceptual degradation in image quality [204].
@@ 11.6 PNG Compression Standard
The Portable Network Graphics (PNG) format is one of the most widely used image formats on the web. Unlike lossy formats like JPEG, PNG provides lossless compression, meaning that no image quality is lost during compression. PNG is commonly used for graphics, logos, icons, and images that require transparency. In this section, we will explore how PNG achieves high-quality compression, including its use of the Deflate compression algorithm, various filter types, and support for alpha channels and interlacing [124].
@@ 11.6.1 Overview of PNG
PNG was developed as a replacement for the older GIF format and provides several advantages, including better compression, support for a broader range of colors, and the inclusion of an alpha channel for transparency. The primary goal of PNG is to achieve efficient, lossless compression without sacrificing image fidelity [18].
@@ 11.6.2 Deflate Compression Algorithm
PNG achieves its compression using the Deflate algorithm, which is a combination of two widely-used techniques: Lempel-Ziv (LZ77) coding and Huffman coding. Deflate is effective in reducing file size while ensuring that no information is lost during compression [140].
Lempel-Ziv (LZ77) Coding: LZ77 works by finding repeated sequences of data in the image file and encoding them as references to earlier occurrences of the same data. This eliminates redundant information, making the file smaller. For example, a pattern that repeats multiple times can be encoded just once, with pointers to each occurrence [116].
Huffman Coding: After LZ77 reduces redundancy, Huffman coding is applied to further compress the data by assigning shorter codes to more frequently occurring symbols and longer codes to less frequent symbols. This variable-length encoding makes the overall file size smaller [128].
@@ 11.6.3 Filter Types in PNG
  • Paeth: This filter predicts the pixel value based on the left, above, and upper-left pixels using the Paeth predictor algorithm, which chooses the most appropriate pixel value to reduce redundancy [166].
@@ 11.6.4 Alpha Channel Support
One of the standout features of PNG is its support for alpha channels, which allows an image to have transparent or semi-transparent pixels. Alpha channels provide additional information about the opacity of each pixel, which is important for graphics that need to blend seamlessly with different backgrounds [171].
@@
[…]
@@ 11.6.5 Interlacing (Adam7 Algorithm)
The Adam7 interlacing algorithm allows a PNG image to load progressively over multiple passes. This is especially useful for large images on the web, where the user can see a rough version of the image before it fully loads, improving the user experience on slow connections [76].
@@ 11.7 Image Compression Quality Assessment
Image compression is an essential process for reducing the size of image files while maintaining acceptable quality. However, evaluating the quality of compressed images requires both objective and subjective methods to ensure that the compression does not degrade the image too much. In this section, we will discuss several techniques for assessing the quality of compressed images, including Peak Signal-to-Noise Ratio (PSNR), Mean Squared Error (MSE), and visual inspection [192].
@@ 11.7.1 Peak Signal-to-Noise Ratio (PSNR)
The Peak Signal-to-Noise Ratio (PSNR) is one of the most commonly used metrics to objectively evaluate the quality of compressed images. PSNR measures the ratio between the maximum possible power of a signal (the original image) and the power of the noise (the compression errors). Higher PSNR values indicate better quality, as the compression introduces less noise [90].
@@ 11.7.2 Mean Squared Error (MSE)
Mean Squared Error (MSE) is a straightforward method for quantifying the difference between the original and compressed images. It calculates the average of the squared differences between the pixel values of the two images [208].
@@
Visual inspection is another important method for assessing image quality. Even if a compressed image has a high PSNR value, it may contain visible artifacts that reduce its subjective quality. Thus, it is often necessary to combine objective measures like PSNR and MSE with subjective visual evaluation to ensure acceptable compression quality [34].
@@
Image recognition is a core task in computer vision, where the goal is to identify objects, patterns, or specific features within an image. The combination of image processing and machine learning allows computers to analyze, interpret, and categorize visual data in various applications, such as object detection, facial recognition, and autonomous driving [141, 161, 51]. In this chapter, we will discuss the fundamental techniques used in image recognition, focusing on feature extraction and pattern recognition methods like support vector machines (SVM) and neural networks.
@@ 12.1 Feature Extraction
Feature extraction is a crucial step in image recognition, as it involves identifying and quantifying meaningful patterns within an image that can be used for classification or other tasks. Features can include texture, color, shape, and spatial relationships between pixels. In this section, we will focus on extracting texture features and shape features [63].
@@ 12.1.1 Texture Features
Texture refers to the visual patterns and surface characteristics in an image, such as smoothness, roughness, and regularity. Texture features can be extracted using several methods, which capture the spatial relationships between pixel intensities and patterns. Two common methods for texture feature extraction are the Gray-Level Co-occurrence Matrix (GLCM) and Local Binary Patterns (LBP) [6].
@@
The GLCM is a matrix that describes how frequently pairs of pixel intensities occur in a specific spatial relationship within an image. It captures the spatial distribution of intensity values and is widely used for texture analysis [41]. The GLCM is computed by considering pairs of pixels separated by a given distance and orientation. Common texture features extracted from GLCM include:
@@
Local Binary Patterns (LBP) is another popular texture feature extraction method that captures the local spatial structure of the image. In LBP, each pixel is compared to its surrounding neighbors, and a binary code is generated based on whether the neighboring pixel values are greater or smaller than the center pixel. This binary pattern is then converted to a decimal value and used as a texture descriptor [150].
@@ 12.1.2 Shape Features
Shape features describe the geometric properties of objects in an image, such as edges, contours, and boundaries. These features are particularly useful for tasks like object recognition and image classification, where the shape of objects is a key characteristic [125].
@@
Edge detection is a fundamental technique for identifying object boundaries by detecting changes in pixel intensity. Common edge detection algorithms include the Sobel, Prewitt, and Canny operators [223].
@@
Contours represent the boundaries of objects in an image and can be used for shape recognition. Contour analysis involves detecting these boundaries and analyzing their geometric properties, such as perimeter, area, and circularity [93].
@@ 12.2 Pattern Recognition Methods
Pattern recognition is the process of classifying data (in this case, image features) into predefined categories. Machine learning algorithms such as Support Vector Machines (SVM) and Neural Networks are commonly used in image recognition tasks to learn patterns and classify images [111].
@@ 12.2.1 Support Vector Machines (SVM)
Support Vector Machines (SVM) are a powerful supervised learning algorithm used for classification tasks. In the context of image recognition, SVMs classify images by finding the optimal hyperplane that separates data points from different classes. The goal of the SVM is to maximize the margin between the hyperplane and the closest data points, known as support vectors [68].
@@
Given a set of labeled training data, an SVM constructs a hyperplane in a high-dimensional space that best separates the classes. If the data is not linearly separable, SVM can use kernel functions (such as the radial basis function, or RBF) to project the data into a higher-dimensional space where it can be separated [105].
@@ 12.2.2 Neural Networks
Neural networks are a class of machine learning models that have become increasingly popular for image recognition tasks. Neural networks consist of layers of interconnected nodes (neurons) that learn to map input features to output labels through a process of training. A simple neural network architecture is the Multilayer Perceptron (MLP), which is made up of an input layer, one or more hidden layers, and an output layer [147].
@@ 12.4 Deep Learning for Image Processing
Deep learning has revolutionized the field of image processing, enabling significant advancements in tasks like image classification, object detection, segmentation, and more. One of the key factors driving this progress is the development of Convolutional Neural Networks (CNNs), which are highly effective for extracting spatial features from images. In this section, we will explore the architecture of CNNs and discuss their applications in tasks such as image classification and object detection [70].
@@ 12.4.1 Convolutional Neural Networks (CNN)
Convolutional Neural Networks (CNNs) are the backbone of most deep learning models for image processing. They are designed to automatically and adaptively learn spatial hierarchies of features from input images. CNNs use a series of convolutional layers to extract features, followed by pooling layers to down-sample the feature maps, and fully connected layers for classification or regression [92].
@@
The typical architecture of a CNN consists of the following components [147]:
@@
6 class SimpleCNN(nn.Module):
@@
9 self.conv1 = nn.Conv2d(1, 16, kernel_size=3, stride=1, padding=1) # Convolutional layer
10 self.conv2 = nn.Conv2d(16, 32, kernel_size=3, stride=1, padding=1) # Convolutional layer
11 self.pool = nn.MaxPool2d(2, 2) # Pooling layer
12 self.fc1 = nn.Linear(32 * 7 * 7, 128) # Fully connected layer
13 self.fc2 = nn.Linear(128, 10) # Output layer for 10 classes

### v2

@@ 11.2.1 Huffman Coding
Huffman coding is a widely used lossless compression algorithm that assigns variable-length codes to symbols (in this case, pixel values), with shorter codes assigned to more frequently occurring symbols. This is based on the observation that, in most images, some pixel values occur more frequently than others. By encoding these frequent values with fewer bits and less frequent values with more bits, Huffman coding achieves compression [moffat2019huffman].
@@ 11.2.2 Run-length Encoding (RLE)
Run-length encoding (RLE) is a simple lossless compression technique that is particularly effective for images with large contiguous regions of the same pixel value, such as binary or grayscale images. The basic idea is to represent consecutive occurrences of the same value as a pair of the value and the count of how many times it repeats [chetan2017performance].
@@ 11.4 Lossy Compression Techniques
Lossy compression techniques are widely used in digital image and multimedia storage because they significantly reduce file sizes by discarding less critical information. Unlike lossless compression, which preserves every detail, lossy compression allows some degradation of image quality in exchange for better compression ratios. The key to effective lossy compression is to remove information that is perceptually less important to the human visual system. In this section, we will cover two of the most important techniques used in lossy compression: the Discrete Cosine Transform (DCT) and the Discrete Wavelet Transform (DWT). These methods are foundational in many compression standards, such as JPEG [koff2006overview].
@@ 11.4.1 Discrete Cosine Transform (DCT)
The Discrete Cosine Transform (DCT) is a fundamental tool in image compression. It converts an image from the spatial domain, where pixel intensities are represented, into the frequency domain, where the image is represented as a sum of cosine functions at various frequencies. This transformation allows the separation of the image into low-frequency components (which represent the bulk of the image’s structure) and high-frequency components (which represent fine details and edges) [ahmed1974discrete].
@@
  • Advantages: The DCT provides good energy compaction and can be efficiently implemented. It is particularly effective for smooth images with few sharp edges [rao2014discrete].
  • Limitations: The DCT assumes that the image is locally smooth, which means it may not perform as well for images with many sharp edges or textures. Additionally, blocking artifacts can occur if the image is divided into small blocks for compression (as in JPEG) [rao2014discrete].
@@ 11.4.2 Discrete Wavelet Transform (DWT)
The Discrete Wavelet Transform (DWT) is another powerful tool used in lossy image compression. Unlike the DCT, which operates on blocks of fixed size, the DWT provides a multi-resolution representation of the image by decomposing it into different levels of detail. The DWT captures both frequency and spatial information, which makes it well-suited for images with varying levels of detail, such as natural scenes [zhang2019wavelet].
@@
The DWT decomposes an image into a set of subbands, each corresponding to different frequency components and spatial resolutions [rozynski2005multi]. Typically, the image is first decomposed into four subbands:
@@ 9 $L$, (LH, HL, $H$) = coeffs2
14 axes[0, 0].imshow($L$, cmap=’gray’)
15 axes[0, 0].set_title(’Approximation ($L$)’)
@@
23 axes[1, 1].imshow($H$, cmap=’gray’)
24 axes[1, 1].set_title(’Diagonal Detail ($H$)’)
@@
  • Advantages: The DWT provides a multi-resolution representation of the image, which allows for efficient compression by discarding or heavily quantizing high-frequency components. It is especially effective for images with detailed textures and varying levels of detail [weeks2003discrete].
  • Limitations: Wavelet-based compression can be more computationally intensive than DCT-based methods. Additionally, wavelet compression can introduce ringing artifacts around sharp edges if the high-frequency components are too aggressively quantized [weeks2003discrete].
@@ 11.5 JPEG Compression Standard
JPEG (Joint Photographic Experts Group) is one of the most commonly used image compression standards, known for its ability to achieve high compression ratios with minimal perceptual loss in image quality. The JPEG algorithm is based on the Discrete Cosine Transform (DCT) and employs several key steps to compress images in a lossy fashion [wallace1990overview]. In this section, we will break down the main steps of the JPEG compression process and explain its widespread use and performance for still image compression.
@@ 11.5.2 Widespread Use and Performance of JPEG
JPEG is widely used for still image compression because it offers a good balance between compression ratio and image quality. It is particularly effective for natural images (e.g., photographs) where slight loss of detail is acceptable. JPEG can achieve compression ratios of 10:1 or higher with little perceptual degradation in image quality [wallace1990overview].
@@ 11.6 PNG Compression Standard
The Portable Network Graphics (PNG) format is one of the most widely used image formats on the web. Unlike lossy formats like JPEG, PNG provides lossless compression, meaning that no image quality is lost during compression. PNG is commonly used for graphics, logos, icons, and images that require transparency. In this section, we will explore how PNG achieves high-quality compression, including its use of the Deflate compression algorithm, various filter types, and support for alpha channels and interlacing [miano1999compressed].
@@ 11.6.1 Overview of PNG
PNG was developed as a replacement for the older GIF format and provides several advantages, including better compression, support for a broader range of colors, and the inclusion of an alpha channel for transparency. The primary goal of PNG is to achieve efficient, lossless compression without sacrificing image fidelity [bourke2001overview].
@@ 11.6.2 Deflate Compression Algorithm
PNG achieves its compression using the Deflate algorithm, which is a combination of two widely-used techniques: Lempel-Ziv (LZ77) coding and Huffman coding. Deflate is effective in reducing file size while ensuring that no information is lost during compression [oswal2016deflate].
Lempel-Ziv (LZ77) Coding: LZ77 works by finding repeated sequences of data in the image file and encoding them as references to earlier occurrences of the same data. This eliminates redundant information, making the file smaller. For example, a pattern that repeats multiple times can be encoded just once, with pointers to each occurrence [louchard1997average].
Huffman Coding: After LZ77 reduces redundancy, Huffman coding is applied to further compress the data by assigning shorter codes to more frequently occurring symbols and longer codes to less frequent symbols. This variable-length encoding makes the overall file size smaller [moffat2019huffman].
@@ 11.6.3 Filter Types in PNG
  • Paeth: This filter predicts the pixel value based on the left, above, and upper-left pixels using the Paeth predictor algorithm, which chooses the most appropriate pixel value to reduce redundancy [roelofs2003png].
@@ 11.6.4 Alpha Channel Support
One of the standout features of PNG is its support for alpha channels, which allows an image to have transparent or semi-transparent pixels. Alpha channels provide additional information about the opacity of each pixel, which is important for graphics that need to blend seamlessly with different backgrounds [rossokhin2014block].
@@
$\text{Pixel }(R=255,G=255,B=255,A=128)\rightarrow\text{50\% transparent white pixel}.$
@@ 11.6.5 Interlacing (Adam7 Algorithm)
The Adam7 interlacing algorithm allows a PNG image to load progressively over multiple passes. This is especially useful for large images on the web, where the user can see a rough version of the image before it fully loads, improving the user experience on slow connections [horak2012influence].
@@ 11.7 Image Compression Quality Assessment
Image compression is an essential process for reducing the size of image files while maintaining acceptable quality. However, evaluating the quality of compressed images requires both objective and subjective methods to ensure that the compression does not degrade the image too much. In this section, we will discuss several techniques for assessing the quality of compressed images, including Peak Signal-to-Noise Ratio (PSNR), Mean Squared Error (MSE), and visual inspection [talukder2010haar].
@@ 11.7.1 Peak Signal-to-Noise Ratio (PSNR)
The Peak Signal-to-Noise Ratio (PSNR) is one of the most commonly used metrics to objectively evaluate the quality of compressed images. PSNR measures the ratio between the maximum possible power of a signal (the original image) and the power of the noise (the compression errors). Higher PSNR values indicate better quality, as the compression introduces less noise [korhonen2012peak].
@@ 11.7.2 Mean Squared Error (MSE)
Mean Squared Error (MSE) is a straightforward method for quantifying the difference between the original and compressed images. It calculates the average of the squared differences between the pixel values of the two images [wang2009mean].
@@
Visual inspection is another important method for assessing image quality. Even if a compressed image has a high PSNR value, it may contain visible artifacts that reduce its subjective quality. Thus, it is often necessary to combine objective measures like PSNR and MSE with subjective visual evaluation to ensure acceptable compression quality [chin1982automated].
@@
Image recognition is a core task in computer vision, where the goal is to identify objects, patterns, or specific features within an image. The combination of image processing and machine learning allows computers to analyze, interpret, and categorize visual data in various applications, such as object detection, facial recognition, and autonomous driving [pak2017review, ren2024deeplearningmachinelearning, feng2024masteringaibigdata]. In this chapter, we will discuss the fundamental techniques used in image recognition, focusing on feature extraction and pattern recognition methods like support vector machines (SVM) and neural networks.
@@ 12.1 Feature Extraction
Feature extraction is a crucial step in image recognition, as it involves identifying and quantifying meaningful patterns within an image that can be used for classification or other tasks. Features can include texture, color, shape, and spatial relationships between pixels. In this section, we will focus on extracting texture features and shape features [guyon2006introduction].
@@ 12.1.1 Texture Features
Texture refers to the visual patterns and surface characteristics in an image, such as smoothness, roughness, and regularity. Texture features can be extracted using several methods, which capture the spatial relationships between pixel intensities and patterns. Two common methods for texture feature extraction are the Gray-Level Co-occurrence Matrix (GLCM) and Local Binary Patterns (LBP) [amadasun1989textural].
@@
The GLCM is a matrix that describes how frequently pairs of pixel intensities occur in a specific spatial relationship within an image. It captures the spatial distribution of intensity values and is widely used for texture analysis [de2013multi]. The GLCM is computed by considering pairs of pixels separated by a given distance and orientation. Common texture features extracted from GLCM include:
@@
Local Binary Patterns (LBP) is another popular texture feature extraction method that captures the local spatial structure of the image. In LBP, each pixel is compared to its surrounding neighbors, and a binary code is generated based on whether the neighboring pixel values are greater or smaller than the center pixel. This binary pattern is then converted to a decimal value and used as a texture descriptor [pietikainen2011computer].
@@ 12.1.2 Shape Features
Shape features describe the geometric properties of objects in an image, such as edges, contours, and boundaries. These features are particularly useful for tasks like object recognition and image classification, where the shape of objects is a key characteristic [mingqiang2008survey].
@@
Edge detection is a fundamental technique for identifying object boundaries by detecting changes in pixel intensity. Common edge detection algorithms include the Sobel, Prewitt, and Canny operators [ziou1998edge].
@@
Contours represent the boundaries of objects in an image and can be used for shape recognition. Contour analysis involves detecting these boundaries and analyzing their geometric properties, such as perimeter, area, and circularity [kukreja2022segmentation].
@@ 12.2 Pattern Recognition Methods
Pattern recognition is the process of classifying data (in this case, image features) into predefined categories. Machine learning algorithms such as Support Vector Machines (SVM) and Neural Networks are commonly used in image recognition tasks to learn patterns and classify images [liu2006pattern].
@@ 12.2.1 Support Vector Machines (SVM)
Support Vector Machines (SVM) are a powerful supervised learning algorithm used for classification tasks. In the context of image recognition, SVMs classify images by finding the optimal hyperplane that separates data points from different classes. The goal of the SVM is to maximize the margin between the hyperplane and the closest data points, known as support vectors [hearst1998support].
@@
Given a set of labeled training data, an SVM constructs a hyperplane in a high-dimensional space that best separates the classes. If the data is not linearly separable, SVM can use kernel functions (such as the radial basis function, or RBF) to project the data into a higher-dimensional space where it can be separated [li2024deeplearningmachinelearning].
@@ 12.2.2 Neural Networks
Neural networks are a class of machine learning models that have become increasingly popular for image recognition tasks. Neural networks consist of layers of interconnected nodes (neurons) that learn to map input features to output labels through a process of training. A simple neural network architecture is the Multilayer Perceptron (MLP), which is made up of an input layer, one or more hidden layers, and an output layer [peng2024deeplearningmachinelearning].
@@ 12.4 Deep Learning for Image Processing
Deep learning has revolutionized the field of image processing, enabling significant advancements in tasks like image classification, object detection, segmentation, and more. One of the key factors driving this progress is the development of Convolutional Neural Networks (CNNs), which are highly effective for extracting spatial features from images. In this section, we will explore the architecture of CNNs and discuss their applications in tasks such as image classification and object detection [hemanth2017deep].
@@ 12.4.1 Convolutional Neural Networks (CNN)
Convolutional Neural Networks (CNNs) are the backbone of most deep learning models for image processing. They are designed to automatically and adaptively learn spatial hierarchies of features from input images. CNNs use a series of convolutional layers to extract features, followed by pooling layers to down-sample the feature maps, and fully connected layers for classification or regression [krizhevsky2012imagenet].
@@
The typical architecture of a CNN consists of the following components [peng2024deeplearningmachinelearning]:
@@
6 class SimpleCNN($n$.Module):
@@
9 self.conv1 = $n$.Conv2d(1, 16, kernel_size=3, stride=1, padding=1) # Convolutional layer
10 self.conv2 = $n$.Conv2d(16, 32, kernel_size=3, stride=1, padding=1) # Convolutional layer
11 self.pool = $n$.MaxPool2d(2, 2) # Pooling layer
12 self.fc1 = $n$.Linear(32 * 7 * 7, 128) # Fully connected layer
13 self.fc2 = $n$.Linear(128, 10) # Output layer for 10 classes